What are the calculation needed to find the answer? Thank you!

Fall-2019 14. Camia Co. contribution margin is $200, after-tax income is $96, and the tax rate is 40%. What are the fixed cos19. Using FIFO process costing, the equivalent units for conversion costs for the month were a. 13,600 b. 14,000 C. 16,000 d.
Answer #1

As per HOMEWORKLIB POLICY we are required to solve the first 4 parts only. For remaining parts you have to post separate question.

Answer 2 Part 14) 3 Fixed Cost = Contribution Margin - (after tax income /(1-tax rate)) = 200 - (96/(1-40%)) 5 = 200 - 160 5

Part 16) Sales = (Fixed cost + Net income before tax)/ contribution margin ratio = ((100000*40%) +(140000/(1-60%)}}/40% = ( 4